OpenStudy (ryan2d):

A pair of linear equations is shown below: y = -x + 1 y = 2x + 4 Which of the following statements best explains the steps to solve the pair of equations graphically?

OpenStudy (ryan2d):

Oh let me fix that

OpenStudy (ryan2d):

a. On a graph, plot the line y = -x + 1 which has y intercept = -1 and slope = 1; and y = 2x + 4 which has y intercept = 2 and slope = 4; and write the coordinates of the point of intersection of the two lines as the solution. b. On a graph, plot the line y = -x + 1 which has y intercept = 1 and slope = 1; and y = 2x + 4 which has y intercept = 1 and slope = 4; and write the coordinates of the point of intersection of the two lines as the solution. c. On a graph, plot the line y = -x + 1 which has y intercept = 1 and slope = -1; and y = 2x + 4 which has y intercept = -2 and slope = 2; and write the coordinates of the point of intersection of the two lines as the solution. d. On a graph, plot the line y = -x + 1 which has y intercept = 1 and slope = -1; and y = 2x + 4 which has y intercept = 4 and slope = 2; and write the coordinates of the point of intersection of the two lines as the solution.

OpenStudy (ryan2d):

Those are the solutions